OpenStudy (anonymous):

A group of friends were voting for their favorite song: two-fifths chose song A, one-sixth of the remaining friends chose song B, and the rest chose song C. If there were thirty friends in all, how many chose song C?

OpenStudy (anonymous):

\[\frac{2}{5}*30\ chose\ song\ A\]Which means 12 people chose song A. ----------------------------------------------------------------- 30-12 or 18 people didn't chose song A. \[\frac{1}{6} * 18\ chose\ song\ B\]Which means 3 people chose song B. ---------------------------------------------------------------- The rest chose C

OpenStudy (texaschic101):

30 people 2/5 chose A -- 2/5 of 30 = 2/5 * 30 = 60/5 = 12 chose A remaining friends (1/6) chose B --- 1/6 * (30 - 12) = 1/6 * 18 = 3 chose B 30 - 12 - 3 = ? chose C